OpenStudy (anonymous):

Which three lengths could be the lengths of the sides of a triangle? 6 cm, 23 cm, 11 cm 10 cm, 15 cm, 24 cm 22 cm, 6 cm, 6 cm 15 cm, 9 cm, 24 cm

OpenStudy (amistre64):

if the shorts add to greater than the long; its good

OpenStudy (amistre64):

6+6 = 12 < 22 .. bad

OpenStudy (amistre64):

11+6 = 17 < 23 ... bad

OpenStudy (anonymous):

cool thank you!!

OpenStudy (amistre64):

10+15 = 25 > 24 ... good

OpenStudy (amistre64):

youre welcome